The ordered pairs (1, 25), (2, 36), (3, 49), (4, 64), and (5, 81) represent a function. What is a rule that represents this function?
Answer by Edwin McCravy(20056) 

( 1, 25 ) = ( 1, 5² ) = ( 1, (1+4)² )
( 2, 36 ) = ( 2, 6² ) = ( 2, (2+4)² )
( 3, 49 ) = ( 3, 7² ) = ( 3, (3+4)² )
( 4, 64 ) = ( 4, 8² ) = ( 4, (4+4)² )
( 5, 81 ) = ( 5, 9² ) = ( 5, (5+4)² )
... ... ...
( N, ? ) = ( N, ? ) = ( N, (N+4)² )
Rule: Start with N, add 4, then square
Edwin
|
|
|